Circular No. 8408

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 NL_{21} (LINEAR) An apparently asteroidal object, designated 2004 NL_{21} on the basis

of LINEAR observations on July 15 and 16 (MPS 110794; discovery observation below), and for which the Minor Planet Center has published several “cometary” orbits (cf. e.g., MPEC 2004-O40, MPO 66393, MPEC 2004-S06), has been reported to be of cometary appearance by R. H. Mc-Naught from observations with the 0.5-m Uppsala Schmidt telescope at Siding Spring. Sixty-second CCD exposures on Sept. 6.6 UT showed a near-asteroidal condensation with a narrow fan tail extending ∼ 40′′ in p.a. 220^{o}; on Sept. 18.5 there was a tail ∼ 60′′ long in p.a. 230^{o}.

2004 UT α_{2000} δ_{2000} Mag. Observer
July 15.34442 23^{h}04^{m}09.^{s}89 −14^{o}10′18′′.4 19.3 LINEAR

The Siding Spring astrometry and other recent observations not already published in the MPS, together with the latest equinox-2000.0 orbital elements (T = 2004 Oct. 13, ω = 18^{o}, Ω = 337^{o}, i = 17^{o}, q = 2.08 AU, P = 7.9 years), residuals, and an ephemeris are given on MPEC 2004-S17.

SUPERNOVAE 2004ec AND 2004ei M. Modjaz, R. Kirshner and P. Challis, Harvard-Smithsonian Center

for Astrophysics; and T. Matheson, National Optical Astronomy Observatory, report that a spectrum (range 350–740 nm) of SN 2004ei (cf. IAUC 8401), obtained by P. Berlind on Sept. 11.43 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), reveals it to be a type-II supernova. The spectrum consists of emission-dominated P-Cyg features of Hα and possibly of Hβ, and a red continuum. Adopting the NED recession velocity of 5755 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hα line is ∼ 6500 km/s. Zero-velocity interstellar Na I D absorption with an equivalent width of ∼ 0.3 nm is detected in the spectrum of SN 2004ei, indicating gas along the line of sight in our galaxy and thus suggesting reddening by dust. Schlegel et al. (1998, Ap.J. 500,

525) estimate E(B − V) = 0.52 mag of Galactic reddening along the line of sight to UGC 2817. A spectrum of SN 2004ec (cf. IAUC 8396), obtained by M. Calkins on Sept. 10.13, shows it to be a type-IIn supernova. The spectrum consists of a blue continuum, indicating young age, and a narrow Hα component of 360 km/s (FWHM) superposed on a weaker and broader component of 1800 km/s (FWHM).

2004 September 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green